150 ml of gas are enclosed in the cylinder under a pressure of 760 torr. What would be the volume be at pressure of 1520 torr?



1
Expert's answer
2021-06-25T06:54:39-0400

.p(1) *V(1)= p(2) *V(2) From which V(2)= p(1) *V(1)/p(2)=150*760/1520=75 ml